### v1

# Hamiltonian Dynamics of Fröhlich Condensates in Classical Systems

Abstract
In the late 1960s, Fröhlich proposed that energy in biological systems may not be entirely dissipated, but stored in an orderly manner, as condensates, to support essential cellular functions. Over 50 years later, evidence for collective protein vibrations in the terahertz (THz) domain seems to confirm these ideas. However, Fröhlich’s model differs remarkably from molecular dynamics (MD), complicating the interpretation of experimental results. As a first step towards linking Fröhlich’s theory with MD-based techniques, Hamiltonian dynamics governing Fröhlich systems are investigated within a classical framework. We show that well-formed condensates can emerge at room temperature from classical Hamiltonians used in normal mode analysis (NMA) to describe protein dynamics. We also suggest a strategy to assess whether standard force fields can capture condensation, paving the way for future studies connecting Fröhlich’s model to biomolecular simulations.
Introduction—Biological systems have always been known for their ability to self-regulate and adapt to external conditions. It is now widely accepted that such properties are the result of a long evolutionary process governed by natural selection that promotes the proper functioning and longevity of organisms [1]. Thus, it is not uncommon to observe biological structures displaying dynamical and energetic properties that appear extraordinary when compared to inert matter. Recently, it was experimentally shown that spectra of proteins like bovine serum albumin (BSA) or R-phycoerythrin (R-PE) are not always thermalized but exhibit sharp sub-THz peaks at high energy input [2, 3]. These results seem consistent with an old theoretical model proposed by H. Fröhlich suggesting that, due to nonlinear processes, energy supplied to a set of THz modes may be specifically channeled into the lowest frequency mode, the so-called condensation effect [4]. Fröhlich condensates are expected to have profound implications for energy storage [4], long-range selective intermolecular forces [3, 5], and cognition [6].
Although Fröhlich’s assumptions are close to the conditions used in the BSA and RP-E experiments, a clear connection with the condensation phenomenon has yet to be established. Recent MD studies have been performed in this regard, but they neither reproduced the observed spectral excitations nor confirmed or refuted Fröhlich’s theory [7, 8]. Limiting factors like the modeling of the energy input or the significant difference between simulation ($\leq 1\ \mu s$) and experimental ($10^{0}-10^{2}\,\mathrm{s}$) timescales might explain these discrepancies. In addition, Fröhlich condensation was originally proposed based on a simplified model, while current MD packages use more refined assumptions [9, 10] involving force fields, thermostats…, which obscures a possible link between the two approaches.
In this work, we show that Fröhlich condensation can be predicted at room temperature by classical Hamiltonian dynamics, at the same level of theory used in MD to model real biomolecules. However, condensation is only possible for specific types of nonlinear coupling with the surrounding environment. Our approach, which incorporates standard thermostats to maintain canonical ensembles, offers a solid basis to highlight the conditions under which a classical system exhibits condensation and verify if standard force fields meet those conditions.
Figure 1: Illustration of a Fröhlich system consisting of a set of oscillators, or modes, in contact with a heat bath at temperature $T$ and a source at temperature $T_{s}$. More details are given in the main text. Figure was inspired by Fig. 1 in [11].
Fröhlich’s model—A typical Fröhlich system is shown in Fig. 1 and consists of 3 components: (1) an ensemble of oscillators, or modes, representing the protein system, (2) a heat bath representing the water/cell environment, and (3) an energy source (e.g., endogenous energy, light…). The protein modes interact linearly with the heat bath and the source at rates $\Phi$ and $\Xi$, respectively, but they are also able to interact with each other. These nonlinear interactions are given by the $\Lambda$ rate and involve pairs of modes interacting via the thermal bath.
Figure 2: Time evolution of harmonic energies in a Fröhlich system made of $9$ protein modes; only the first 5 modes are shown. Results were obtained by solving Eqs (1) numerically with frequencies ranging from $\omega_{1}=0.2\ \text{THz}$ to $\omega_{9}=1\ \text{THz}$ with a $0.1$-THz increment [[Whilefrequenciesaretypicaloflow-frequencymodesinproteins(seeforexample][), rateconstantvalueswereessentiallychosentoshowcondensationwithinatimeframeaccessibletoMDsimulations.]go1983dynamics]. Other parameters: $\Phi_{i}=5\cdot 10^{-5}\,\text{ps}^{-1}$, $\Lambda_{ij}=5\cdot 10^{-5}\,\text{ps}^{-1}$, $T=300\ \text{K}$, $T_{s}=3000\ \text{K}$. $\Xi_{i}$ was set to $5\cdot 10^{-6}\,\text{ps}^{-1}$ at $50$ ns ($\Xi_{i}=0$ before $50$ ns).
Based on the above assumptions, Fröhlich suggested a set of equations referred here as Fröhlich rate equations (FRE) to describe the energy of each protein mode. Calling $\langle J_{i}\rangle$ the average action of mode with frequency $\omega_{i}$, the FRE are given in the classical limit by [3, 13]:
$\begin{split}\dot{\langle J_{i}\rangle}&=\Phi_{i}\left(\frac{kT}{\omega_{i}}-\langle J_{i}\rangle\right)+\Xi_{i}\left(\frac{kT_{s}}{\omega_{i}}-\langle J_{i}\rangle\right)+\\ &\qquad\sum\limits_{j=1}^{N}\Lambda_{ij}\left(\langle J_{j}\rangle-\langle J_{i}\rangle+\frac{\omega_{j}-\omega_{i}}{kT}\langle J_{i}\rangle\langle J_{j}\rangle\right)\\ &\qquad\qquad\qquad\qquad\mathrm{with}\ i=1\dots N,\end{split}$ (1)
where $N$ is the number of modes and the RHS includes the 3 coupling types introduced above: $\Phi$ and $\Xi$ (linear) and $\Lambda$ (nonlinear). Remarkably, each term was originally postulated from the condition that energies $E_{i}=\omega_{i}\langle J_{i}\rangle$ are always equal to $kT$ or $kT_{s}$ in the stationary state, depending on whether the bath or the source is involved [4]. For instance, if only the first term in the RHS is considered, $\dot{\langle J_{i}\rangle}=0$ will give $\omega_{i}\langle J_{i}\rangle=kT$ for all $i$.
[…]
Hamiltonian dynamics—A class of quantum Hamiltonians was proposed by Wu and Austin [14, 15] to describe the microscopic dynamics of Fröhlich systems. In this formalism, both the heat bath and the source are modelled as two additional sets of modes interacting with the protein. We focus here on the classical version of these Hamiltonians, given by $H=H_{0}+H_{int}$ where
$\begin{split}H_{0}=\sum\limits_{i=1}^{N}&\frac{p_{i}^{2}}{2m_{i}}+\frac{1}{2}m_{i}\omega_{i}^{2}q_{i}^{2}\>+\\ &\quad\sum\limits_{k=1}^{N_{B}}\frac{p_{k}^{{\scriptscriptstyle(B)}2}}{2m_{k}^{\scriptscriptstyle{(B)}}}+\frac{1}{2}m_{k}^{\scriptscriptstyle{(B)}}{\omega_{k}^{{\scriptscriptstyle(B)}2}}q_{k}^{{\scriptscriptstyle(B)}2}\>+\\ &\qquad\qquad\sum\limits_{l=1}^{N_{S}}\frac{p_{l}^{{\scriptscriptstyle(S)}2}}{2m_{l}^{\scriptscriptstyle{(S)}}}+\frac{1}{2}m_{l}^{\scriptscriptstyle(S)}{\omega_{l}^{{\scriptscriptstyle(S)}2}}q_{l}^{{\scriptscriptstyle(S)}2},\end{split}$ (2)
and $H_{int}$ is the interaction Hamiltonian such that
$H_{int}=\sum\limits_{ik}\phi_{ik}q_{i}q_{k}^{{\scriptscriptstyle(B)}}+\sum\limits_{ik}\xi_{il}q_{i}q_{l}^{{\scriptscriptstyle(S)}}+\sum\limits_{ijk}\lambda_{ijk}q_{i}q_{j}q_{k}^{{\scriptscriptstyle(B)}}.$ (3)
Here $p_{i}$, $p_{k}^{\scriptscriptstyle(B)}$, $p_{l}^{\scriptscriptstyle(S)}$ and $q_{i}$, $q_{k}^{\scriptscriptstyle(B)}$, $q_{l}^{\scriptscriptstyle(S)}$ are the impulsions and the positions of the protein modes, the bath and the source, respectively, while $m_{i}$, $m_{k}^{\scriptscriptstyle{(B)}}$, $m_{l}^{\scriptscriptstyle{(S)}}$ and $\omega_{i}$, $\omega_{k}^{\scriptscriptstyle{(B)}}$, $\omega_{l}^{\scriptscriptstyle{(S)}}$ are their associated masses and frequencies. $N$, $N_{B}$ and $N_{S}$ are the numbers of modes in each set. Finally, $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ are the coupling coefficients related to the 3 types of interaction originally introduced by Fröhlich.
Previously, we have shown how the FRE can be recovered from a Hamiltonian similar to the one above [13]. This derivation supposes that the heat bath and the source are kept at temperature $T$ and $T_{s}$, respectively, and that nonlinear interactions are primarily driven by low-frequency modes of the bath. The latter assumption implies that resonances of the type $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle{(B)}}=0$ are neglegible over resonances of the type $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle{(B)}}=0$. Moreover, most derivations of the FRE are based on perturbation theory assuming high-order perturbative terms can either be included in the rate constants [15], or are negligible [16]. This contrasts with many studies on rate equations showing both quantitative and qualitative changes when higher-order terms are considered [17, 18]. When using perturbation theory up to second order, rate constants are related to $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ as follows [13]
$\Phi_{i}=\sum_{k}\frac{\alpha\,\phi_{ik}^{2}}{m_{i}m_{k}^{\scriptscriptstyle{(B)}}\omega_{i}^{2}}\,\delta(\omega_{i}-\omega_{k}^{\scriptscriptstyle(B)})$ (4a)
$\Xi_{i}=\sum_{l}\frac{\alpha\,\xi_{il}^{2}}{m_{i}m_{l}^{\scriptscriptstyle{(S)}}\omega_{i}^{2}}\,\delta(\omega_{i}-\omega_{l}^{\scriptscriptstyle(S)})$ (4b)
$\Lambda_{ij}=\sum_{k}\frac{2\alpha\,\lambda_{ijk}^{2}}{m_{i}m_{j}m_{k}^{\scriptscriptstyle{(B)}}}\frac{kT\delta(\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)})}{\omega_{i}\omega_{j}(\omega_{j}-\omega_{i})^{2}}$ (4c)
where $\alpha$ is a time scaling factor to be determined. Note that $\Phi_{i}$, $\Xi_{i}$ and $\Lambda_{ij}$ depend on resonance conditions meaning that off-resonance interactions are supposed to have little impact on the dynamics.
Figure 3: Time evolution of the total energies, i.e., the sum of kinetic and potential energies, in a Fröhlich system made of $9$ protein modes; only the first 5 modes are shown. Energies were computed as moving averages over $300$ ns. Results were obtained from Hamiltonian dynamics by keeping the bath at $T=300\ \text{K}$ and the source at $T_{s}=3000\ \text{K}$. Protein frequencies were set from $\omega_{1}=0.2$ to $\omega_{9}=1$ THz with $0.1$-THz increment. $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ were taken from Eqs. (5) and (6) with $\phi=1.0$ and $\xi=0.4$ (from $100$ ns). Top: $\lambda=0.0$, bottom: $\lambda=0.95\ \text{K}^{-1/2}$. Masses were all set to unity. Curves in black correspond to the predictions of the FRE using Eqs. (4) with $\alpha=0.02$ ps; the solid black line shows mode 1 while dashed curves correspond to secondary modes.
Numerical simulations—To check the validity of the FRE and of the condensation effect, Hamilton’s equations were integrated from the sum of Hamiltonians (2) and (3). A velocity Verlet integration scheme was applied. The bath oscillators were kept at $T=300\,\text{K}$ by coupling them to a Langevin thermostat while another Langevin thermostat was used to maintain the source oscillators at $T_{s}$, here treated as a free parameter. Each thermostat was tested individually for the bath and the source. This is shown in Fig. S2, where each set exhibits the right temperature as computed from ensemble and time averages.
To get closer to Fröhlich’s settings and limit the number of calculations of pairwise and triplet interactions, simulations were run by applying strict resonances only, i.e., coefficients $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ were all set to zero except at resonance. Coupling coefficients were also supposed to be proportional to the square root of the masses and to the frequencies. For instance, in the case of $\phi$ coupling, the following coefficients were used:
$\phi_{ik}=\left\{\begin{array}[]{l}\phi\,\sqrt{m_{i}m_{k}^{\scriptscriptstyle{(B)}}}\omega_{i}\omega_{k}^{\scriptscriptstyle(B)}/N_{\phi},\>\text{when}\ \omega_{i}=\omega_{k}^{\scriptscriptstyle(B)}\vspace{0.4cm}\\ 0\>\ \text{otherwise,}\end{array}\right.$ (5)
where $\phi$ is a unitless parameter and $N_{\phi}$ is the number of resonances related to $\phi$ interactions. A similar expression was used for the $\xi_{il}$ coefficients by introducing the parameter $\xi$ and the number $N_{\xi}$ of $\xi$-resonances (not shown). Similarly, coefficients related to $\lambda$-coupling were set as
$\lambda_{ijk}=\left\{\begin{array}[]{l}\lambda\,\sqrt{m_{i}m_{j}m_{k}^{\scriptscriptstyle{(B)}}}\omega_{i}\omega_{j}\omega_{k}^{\scriptscriptstyle(B)}/N_{\lambda},\vspace{0.1cm}\\ \qquad\qquad\qquad\text{when}\ \omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0\vspace{0.3cm}\\ 0\>\ \text{otherwise}\end{array}\right.$ (6)
where $\lambda$ is given in $\text{K}^{-1/2}$ units. Assuming the bath and the source are much bigger entities than the protein, their frequencies spectra were always set to generate the maximum number of resonances possible, i.e., $N_{\phi}=N$, $N_{\xi}=N$ and $N_{\lambda}=N(N-1)$ where $N$ is the number of protein modes. Physical intuition for setting $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ from Eqs. (5) and (6) is given in the next sections.
Results of our simulations are displayed in Fig. 3, where we observed that the two main features of Fröhlich systems could be reproduced, that is, energy equipartition when nonlinear interactions are turned off ($\lambda=0$), and condensation in mode 1 at high $\lambda$ value. Although the FRE show reasonable agreement with real dynamics, they also tend to miscalculate the energy available in condensates, either by overestimating (Fig. 3 bottom) or underestimating it (Fig. S3), showing the limits of Eqs. (1) combined with Eq. (4). Dynamics also revealed the formation of strong condensates in specific regions of the parameter space characterized by reasonable source temperature and energies of a dozen of $kT$ (Fig. S3), consistent with typical energy values required to induce natural conformational changes in proteins. This observation goes against a previous numerical study by Reimers et al. affirming that strong condensates are unlikely to happen in a biological environment due to the high-energy values they require [11].
Figure 4: Condensation index $\eta$ of a Fröhlich system as deduced from Hamiltonian dynamics for different $\lambda$ values. $\lambda_{ijk}$ coefficients were calculated from Eqs. (6) under different types of resonance: $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ (Fröhlich), $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$ (Lifshits) and $\omega_{i}\pm\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ (combination of both). Other parameters remain the same as Fig. 3.
Lifshits terms—We mentioned above that the FRE implicitly assume that nonlinear resonances of the type $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ are predominant over other types of resonance like $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$. Lifshits [19] identified the latter type as a possible hindrance to condensation. To investigate this, another round of simulations was run by computing the $\lambda_{ijk}$ coefficients from the same expression given in Eq. (6) but setting them to be nonzero only when $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$. Extra simulations were performed in the case $\omega_{i}\pm\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$, i.e., when both Fröhlich and Lifshits resonances were involved.
Results are depicted in Fig. 4 highlighting the condensation index $\eta$ as a function of the $\lambda$ parameter. Here $\eta$ was defined as the energy of mode 1 in the stationary state divided by the average energy over the secondary modes, i.e., $\eta=E_{1}/\langle E_{i}\rangle_{i=2..N}$. Thus, $\eta=1$ indicates energy equipartition whereas $\eta>1$ implies condensation. From Fig. 4, we see that resonances involving high-frequency modes of the bath tend to destroy the condensation effect, regardless of whether Fröhlich resonances are included or not. Although it was suggested that the magnitude of Lifshits resonances is likely to be negligible in practice [16], investigation of Fröhlich condensates in real structures would require careful evaluation of these contributions as deduced from all-atom force fields.
Protein-bath vs protein-protein coupling—In addition to nonlinear resonances, we explored the importance of the heat bath in nonlinear coupling. To this purpose, we modified the cubic potential in Eq. (3), replacing the $\lambda_{ijk}q_{i}q_{j}q_{k}^{{\scriptscriptstyle(B)}}$ term with a $\lambda_{ijk}q_{i}q_{j}q_{k}$ term involving only protein modes. This modification was sufficient to completely suppress Fröhlich condensation across all tested parameter regions where the phenomenon was originally observed. In all cases, energy equipartition was observed (not shown). While we cannot rule out the existence of Fröhlich condensates in unexplored regions of the parameter space, our results strongly indicate that bath-mediated coupling is essential for inducing condensation.
Constant coupling—In their study, Reimers et al. [11] reported that they had investigated Fröhlich condensation over a wide region of the parameter space of the quantum Wu-Austin Hamiltonian, concluding with the non-existence of strong and coherent condensates in living matter. However, only constant coupling parameters were considered in their Hamiltonian, drastically limiting the exploration of the condensates. As mentioned above, our simulations were run by setting $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ proportional to the square root of the masses and to the frequencies. Such a choice was motivated by the fact that, at thermal equilibrium, averaged potential energies $\frac{1}{2}m_{i}\omega_{i}^{2}\langle q_{i}^{2}\rangle$ should equal $kT/2$ or $kT_{s}/2$ for the bath or the energy source, respectively. Thus, 2 modes with the same mass but different frequencies, say $\omega_{1}<\omega_{2}$, should satisfy $\langle q_{1}^{2}\rangle>\langle q_{2}^{2}\rangle$ in order to preserve equilibrium condition, meaning that low-frequency modes will have naturally higher amplitude than high-frequency ones. In that case, an interaction potential with constant coefficients (e.g., $\phi_{ik}q_{i}q_{k}^{{\scriptscriptstyle(B)}}$ with $\phi_{ik}=\phi$) will generate stronger interactions between low-frequency modes. This observation can also be made for masses, i.e., modes with a large mass will have a stronger impact on the interaction energy. Setting coefficients as in Eqs (5) and (6) enables to circumvent this issue, ensuring that all the modes contribute equally to the potential.
Despite the above, the impact of a constant coupling on a Fröhlich system was still investigated. In this scenario, our results showed that the interaction energy decreases exponentially as Fröhlich condensation takes place. More precisely, the interaction energy in the lowest frequency mode becomes of the same order of magnitude as the harmonic energy, then rapidly diverges into negative values (not shown). This result is clearly not compatible with our Hamiltonian model whereby $H_{int}$ is always considered a small perturbation as compared to $H_{0}$. Again, using Eqs (5) and (6) helped fix this issue by keeping the interaction energy reasonably low even when strong condensation occurs (Fig. S4, see also note [20]).
Discussion and perspectives—The present study deals with the foundations of the condensation phenomenon proposed by Fröhlich in the 1960s aimed at explaining coherent excitations in biomolecular systems [4]. Motivated by recent experimental findings [2, 3], we simulated the dynamics of a general classical Hamiltonian that incorporates all the key elements of Fröhlich’s models. Our results show that condensation, including strong condensates with reasonable energy values, can emerge under specific types of nonlinear coupling, supporting the physical plausibility of Fröhlich’s effect. However, small deviations from the original model, for example, via the addition of resonances involving high-frequency modes of the bath or the use of a nonlinear potential without bath mediation, result in suppressing the phenomenon. These factors should be carefully examined when investigating the condensation effect in real biomolecular structures.
Given the large number of free parameters involved in the dynamics of Fröhlich’s systems, a full exploration of the parameter space is beyond the scope of this letter. Although off-resonance coupling [21] or higher-order coupling mechanisms could bring more realistic characteristics to the model [22], a more practical approach would be to tune coupling parameters directly from standard MD force fields. Possible strategies include normal mode analysis (NMA) to generate the set of eigenvectors needed to transition from real space to the space of protein normal modes [23, 24]. In this case, not only normal frequencies of the protein and the heat bath could be deduced but eigenvectors may also be used to evaluate the force field function in the space of modes. This would allow coupling coefficients in Eq. (3) to be deduced and used in our model to verify whether condensation can indeed arise. Finally, such an approach should give further insights into the reasons why previous MD studies have been unsuccessful in detecting Fröhlich’s effect and if missing ingredients are required.
Acknowledgments—This work and J. Preto’s postdoctoral fellowship were supported by the European Union’s Horizon Research and Innovation Programme under Grant Agreement No. 964203 (FET-Open LINkS project).

### v2

@@
-Hamiltonian Dynamics of Fröhlich Condensates in Classical Systems
+Hamiltonian Model for Energy Condensation in Classical Systems: Relevance to Proteins
 Abstract
-In the late 1960s, Fröhlich proposed that energy in biological systems may not be entirely dissipated, but stored in an orderly manner, as condensates, to support essential cellular functions.
-Over 50 years later, evidence for collective protein vibrations in the terahertz (THz) domain seems to confirm these ideas.
-However, Fröhlich’s model differs remarkably from molecular dynamics (MD), complicating the interpretation of experimental results.
-As a first step towards linking Fröhlich’s theory with MD-based techniques, Hamiltonian dynamics governing Fröhlich systems are investigated within a classical framework.
-We show that well-formed condensates can emerge at room temperature from classical Hamiltonians used in normal mode analysis (NMA) to describe protein dynamics.
-We also suggest a strategy to assess whether standard force fields can capture condensation, paving the way for future studies connecting Fröhlich’s model to biomolecular simulations.
+Recent experimental evidence for collective protein vibrations in the terahertz (THz) domain indicates that energy in biomolecular systems can self-organize in an orderly manner, as anticipated by Fröhlich’s theory of condensates within a quantum framework.
+As a first step to bridge THz experiments with theory, we study the Hamiltonian dynamics of a classical network of coupled normal modes representing Fröhlich-type systems.
+Our results demonstrate that biologically relevant condensates can emerge at room temperature under appropriate nonlinear coupling schemes.
+The condensation mechanism remains robust also when the original Fröhlich resonance conditions are relaxed.
 Introduction—Biological systems have always been known for their ability to self-regulate and adapt to external conditions.
 It is now widely accepted that such properties are the result of a long evolutionary process governed by natural selection that promotes the proper functioning and longevity of organisms [1].
-Thus, it is not uncommon to observe biological structures displaying dynamical and energetic properties that appear extraordinary when compared to inert matter.
-Recently, it was experimentally shown that spectra of proteins like bovine serum albumin (BSA) or R-phycoerythrin (R-PE) are not always thermalized but exhibit sharp sub-THz peaks at high energy input [2, 3].
+As a result, biological structures often display energetic properties that appear extraordinary when compared to inert matter.
+Recent experiments have shown that spectra of proteins like bovine serum albumin (BSA) or R-phycoerythrin (R-PE) are not always thermalized but exhibit sharp sub-THz peaks upon laser light excitation [2, 3].
+In the case of R-PE, such low-frequency modes could be triggered not only upon (local) electronic excitation of protein fluorochromes, but also upon a temperature increase [4].
+This proves that a disordered and delocalized form of energy like heat, which should simply distribute among all the normal modes of the protein, is instead driven into a well-defined collective oscillation.
 These results seem consistent with an old theoretical model proposed by H.
-Fröhlich suggesting that, due to nonlinear processes, energy supplied to a set of THz modes may be specifically channeled into the lowest frequency mode, the so-called condensation effect [4].
-Fröhlich condensates are expected to have profound implications for energy storage [4], long-range selective intermolecular forces [3, 5], and cognition [6].
-Although Fröhlich’s assumptions are close to the conditions used in the BSA and RP-E experiments, a clear connection with the condensation phenomenon has yet to be established.
-Recent MD studies have been performed in this regard, but they neither reproduced the observed spectral excitations nor confirmed or refuted Fröhlich’s theory [7, 8].
-Limiting factors like the modeling of the energy input or the significant difference between simulation ($\leq 1\ \mu s$) and experimental ($10^{0}-10^{2}\,\mathrm{s}$) timescales might explain these discrepancies.
-In addition, Fröhlich condensation was originally proposed based on a simplified model, while current MD packages use more refined assumptions [9, 10] involving force fields, thermostats…, which obscures a possible link between the two approaches.
-In this work, we show that Fröhlich condensation can be predicted at room temperature by classical Hamiltonian dynamics, at the same level of theory used in MD to model real biomolecules.
-However, condensation is only possible for specific types of nonlinear coupling with the surrounding environment.
-Our approach, which incorporates standard thermostats to maintain canonical ensembles, offers a solid basis to highlight the conditions under which a classical system exhibits condensation and verify if standard force fields meet those conditions.
+Fröhlich [5], suggesting that, due to nonlinear processes, energy supplied to a set of THz modes may be specifically channeled into the lowest frequency mode, the so-called energy condensation.
+This effect could have profound implications for energy storage [5], long-range selective intermolecular forces [4, 3, 6], and cognition [7].
+Starting from the numerical integration of Fröhlich’s rate equations (FREs) in the semiclassical and classical limits, Reimers et al.
+[8] identified regions of the parameter space leading to ”weak” and ”strong” condensates, which were discussed with respect to their possible biological relevance.
+The authors have also explored the quantum coherence of condensates by mapping the rate equations into a quantum Hamiltonian framework.
+Experimental results mentioned above have also motivated force-field-based simulations, but the latter neither reproduced the observed spectral excitations nor confirmed or refuted Fröhlich’s theory [9, 10].
+Possible reasons for these discrepancies include the modeling of the energy input as well as the significant difference between simulation ($\leq 1\ \mu s$) and experimental ($10^{0}-10^{2}\,\mathrm{s}$) timescales.
+In this work, we introduce a generic model showing that energy condensation of Fröhlich type can occur from simple classical Hamiltonian dynamics.
+Using a system of coupled oscillators representing relevant protein modes, the nearby environment (bath) and an energy source, we demonstrate that condensation in the lowest-frequency mode of the protein emerges for specific types of nonlinear (order-three) coupling with the environment.
+Energy injection and dissipation is achieved by coupling one thermostat to the bath and another to the source at a higher temperature placing the protein in out-of-equilibrium conditions.
+With this setup, we found that energy condensation persisted when coupling constants and/or friction coefficients were rescaled over two orders of magnitude, highlighting the robustness of the phenomenon.
+While the model does not aim at explaining the experimental results of specific proteins, it establishes the foundations to explore Fröhlich condensation in quasi-classical open systems, as biomolecules are generally modeled.
+Furthermore, projection of force field-based trajectories in the normal mode space is routinely performed to dissect the atomistic contributions to the essential functional dynamics of proteins (mainly driven by low-frequency modes) [11].
+Using this approach, out-of-equilibrium force field-based trajectories of proteins subjected to an instantaneous external perturbation have been well described in terms of coupled normal modes-models [12], which have highlighted the relevant role of order-three coupling in the energy redistribution among protein modes.
+In this paper, we first formulate the problem using Fröhlich’s rate equations in the classical limit.
+We then introduce the classical Hamiltonian model and numerically examine the emergence of condensates across different parameter sets and nonlinear coupling conditions.
 Figure 1: Illustration of a Fröhlich system consisting of a set of oscillators, or modes, in contact with a heat bath at temperature $T$ and a source at temperature $T_{S}$.
@@
 1 in [8].
-Fröhlich’s model—A typical Fröhlich system is shown in Fig.
-1 and consists of 3 components: (1) an ensemble of oscillators, or modes, representing the protein system, (2) a heat bath representing the water/cell environment, and (3) an energy source (e.g., endogenous energy, light…).
+Fröhlich systems—A typical Fröhlich system is shown in Fig.
+1 and consists of 3 components: (1) an ensemble of oscillators, or modes, representing the protein of interest, (2) a heat bath representing the water/cell environment, and (3) an energy source (e.g., endogenous energy, light…).
 The protein modes interact linearly with the heat bath and the source at rates $\Phi$ and $\Xi$, respectively, but they are also able to interact with each other.
@@
 Figure 2: Time evolution of harmonic energies in a Fröhlich system made of $9$ protein modes; only the first 5 modes are shown.
-Results were obtained by solving Eqs (1) numerically with frequencies ranging from $\omega_{1}=0.2\ \text{THz}$ to $\omega_{9}=1\ \text{THz}$ with a $0.1$-THz increment [[Whilefrequenciesaretypicaloflow-frequencymodesinproteins(seeforexample][), rateconstantvalueswereessentiallychosentoshowcondensationwithinatimeframeaccessibletoMDsimulations.]go1983dynamics].
-Other parameters: $\Phi_{i}=5\cdot 10^{-5}\,\text{ps}^{-1}$, $\Lambda_{ij}=5\cdot 10^{-5}\,\text{ps}^{-1}$, $T=300\ \text{K}$, $T_{s}=3000\ \text{K}$.
+Results were obtained by solving Eqs (1) numerically with frequencies uniformly distributed ranging from $\omega_{1}=0.2\ \text{THz}$ to $\omega_{9}=1\ \text{THz}$ with a $0.1$-THz increment [[Whilefrequenciesaretypicaloflow-frequencymodesinproteins(seeforexample][), rateconstantvalueswereessentiallychosentoshowcondensationwithinatimeframeaccessibletoMDsimulations.]go1983dynamics].
+Other parameters are $\Phi_{i}=5\cdot 10^{-5}\,\text{ps}^{-1}$, $\Lambda_{ij}=5\cdot 10^{-5}\,\text{ps}^{-1}$, T=300KT=300\ $\text${K}, $T_{S}=3000\ \text{K}$.
 $\Xi_{i}$ was set to $5\cdot 10^{-6}\,\text{ps}^{-1}$ at $50$ ns ($\Xi_{i}=0$ before $50$ ns).
-Based on the above assumptions, Fröhlich suggested a set of equations referred here as Fröhlich rate equations (FRE) to describe the energy of each protein mode.
-Calling $\langle J_{i}\rangle$ the average action of mode with frequency $\omega_{i}$, the FRE are given in the classical limit by [3, 13]:
+Based on the above assumptions, Fröhlich suggested a set of rate equations (FREs) to describe the energy of each protein mode.
+Calling $\langle J_{i}\rangle$ the average action of mode $i$ with frequency $\omega_{i}$, the FREs are given in the classical limit by [3, 14]:
 $\begin{split}\dot{\langle J_{i}\rangle}&=\Phi_{i}\left(\frac{kT}{\omega_{i}}-\langle J_{i}\rangle\right)+\Xi_{i}\left(\frac{kT_{S}}{\omega_{i}}-\langle J_{i}\rangle\right)+\\ &\qquad\sum\limits_{j=1}^{N}\Lambda_{ij}\left(\langle J_{j}\rangle-\langle J_{i}\rangle+\frac{\omega_{j}-\omega_{i}}{kT}\langle J_{i}\rangle\langle J_{j}\rangle\right)\\ &\qquad\qquad\qquad\qquad\mathrm{with}\ i=1\dots N,\end{split}$ (1)
@@
 For instance, if only the first term in the RHS is considered, $\dot{\langle J_{i}\rangle}=0$ will give $\omega_{i}\langle J_{i}\rangle=kT$ for all $i$.
-[…]
+The relationship between our rate constants and those of the original Fröhlich’s equations is given in the Supplemental Material [15].
+Two important properties of Fröhlich systems can be deduced from the FREs.
+First, switching off nonlinear interactions, i.e., $\Lambda_{ij}=0$, always leads to energy equipartition regardless of whether the source is active or not.
+This is illustrated in Fig.
+S1 in [15], where the energy source is activated at 50ns50\ $\text${ns}.
+Secondly, if $\Lambda_{ij}$ is sufficiently large, energy will be channeled into the lowest frequency mode, provided that energy is supplied at a high enough rate $\Xi_{i}$.
+This phenomenon, known as Fröhlich condensation, is depicted in Fig.
+2 and was originally suggested to explain the emergence of specific low-frequency modes in biomolecular structures.
 Hamiltonian dynamics—A class of quantum Hamiltonians was proposed by Wu and Austin [16, 17] to describe the microscopic dynamics of Fröhlich systems.
@@
 $H_{int}=\sum\limits_{ik}\phi_{ik}q_{i}q_{k}^{{\scriptscriptstyle(B)}}+\sum\limits_{ik}\xi_{il}q_{i}q_{l}^{{\scriptscriptstyle(S)}}+\sum\limits_{ijk}\lambda_{ijk}q_{i}q_{j}q_{k}^{{\scriptscriptstyle(B)}}.$ (3)
-Here $p_{i}$, $p_{k}^{\scriptscriptstyle(B)}$, $p_{l}^{\scriptscriptstyle(S)}$ and $q_{i}$, $q_{k}^{\scriptscriptstyle(B)}$, $q_{l}^{\scriptscriptstyle(S)}$ are the impulsions and the positions of the protein modes, the bath and the source, respectively, while $m_{i}$, $m_{k}^{\scriptscriptstyle{(B)}}$, $m_{l}^{\scriptscriptstyle{(S)}}$ and $\omega_{i}$, $\omega_{k}^{\scriptscriptstyle{(B)}}$, $\omega_{l}^{\scriptscriptstyle{(S)}}$ are their associated masses and frequencies.
+Here $p_{i}$, $p_{k}^{\scriptscriptstyle(B)}$, $p_{l}^{\scriptscriptstyle(S)}$ and $q_{i}$, $q_{k}^{\scriptscriptstyle(B)}$, $q_{l}^{\scriptscriptstyle(S)}$ are the momenta and the positions of the protein modes, the bath and the source, respectively, while $m_{i}$, $m_{k}^{\scriptscriptstyle{(B)}}$, $m_{l}^{\scriptscriptstyle{(S)}}$ and $\omega_{i}$, $\omega_{k}^{\scriptscriptstyle{(B)}}$, $\omega_{l}^{\scriptscriptstyle{(S)}}$ are their associated masses and frequencies.
 $N$, $N_{B}$ and $N_{S}$ are the numbers of modes in each set.
 Finally, $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ are the coupling coefficients related to the 3 types of interaction originally introduced by Fröhlich.
-Previously, we have shown how the FRE can be recovered from a Hamiltonian similar to the one above [13].
+Previously, we have shown how the FREs can be recovered from a Hamiltonian similar to the one above [14].
 This derivation supposes that the heat bath and the source are kept at temperature $T$ and $T_{S}$, respectively, and that nonlinear interactions are primarily driven by low-frequency modes of the bath.
-The latter assumption implies that resonances of the type $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle{(B)}}=0$ are neglegible over resonances of the type $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle{(B)}}=0$.
-Moreover, most derivations of the FRE are based on perturbation theory assuming high-order perturbative terms can either be included in the rate constants [15], or are negligible [16].
+The latter assumption implies that resonances of the type $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle{(B)}}=0$ are negligible over resonances of the type $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle{(B)}}=0$.
+Moreover, most derivations of the FREs are based on perturbation theory assuming high-order perturbative terms can either be included in the rate constants [17], or are negligible [18].
 This contrasts with many studies on rate equations showing both quantitative and qualitative changes when higher-order terms are considered [19, 20].
@@
 Energies were computed as moving averages over $300$ ns.
-Results were obtained from Hamiltonian dynamics by keeping the bath at $T=300\ \text{K}$ and the source at $T_{s}=3000\ \text{K}$.
-Protein frequencies were set from $\omega_{1}=0.2$ to $\omega_{9}=1$ THz with $0.1$-THz increment.
+Results were obtained from Hamiltonian dynamics by keeping the bath at T=300KT=300\ $\text${K} and the source at $T_{S}=3000\ \text{K}$.
+Protein frequencies were set uniformly from $\omega_{1}=0.2$ to $\omega_{9}=1$ THz with $0.1$-THz increment.
 $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ were taken from Eqs.
-(5) and (6) with $\phi=1.0$ and $\xi=0.4$ (from $100$ ns).
+(5) with $\phi=1.0$ and $\xi=0.4$ (from $100$ ns).
 Top: $\lambda=0.0$, bottom: $\lambda=0.95\ \text{K}^{-1/2}$.
 Masses were all set to unity.
-Curves in black correspond to the predictions of the FRE using Eqs.
+Friction coefficients for the bath and source thermostats were both set to 0.1 $\mathrm{ps}^{-1}$.
+Curves in black correspond to the predictions of the FREs using Eqs.
 (4) with $\alpha=0.02$ ps; the solid black line shows mode 1 while dashed curves correspond to secondary modes.
-Numerical simulations—To check the validity of the FRE and of the condensation effect, Hamilton’s equations were integrated from the sum of Hamiltonians (2) and (3).
+Predictions of FREs are similar to those expected based on ref.
+[8]
+Numerical simulations—To check the validity of the FREs and of the condensation effect, Hamilton’s equations were integrated from $H_{0}+H_{int}$ given by (2) and (3).
 A velocity Verlet integration scheme was applied.
 The bath oscillators were kept at $T=300\,\text{K}$ by coupling them to a Langevin thermostat while another Langevin thermostat was used to maintain the source oscillators at $T_{S}$, here treated as a free parameter.
-Each thermostat was tested individually for the bath and the source.
+Bath and source thermostats were tested individually.
 This is shown in Fig.
-S2, where each set exhibits the right temperature as computed from ensemble and time averages.
-To get closer to Fröhlich’s settings and limit the number of calculations of pairwise and triplet interactions, simulations were run by applying strict resonances only, i.e., coefficients $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ were all set to zero except at resonance.
+S2 in [15], where each set exhibits the right temperature as computed from ensemble and time averages.
+To get closer to Fröhlich’s settings, simulations were run by applying strict resonances only, i.e., coefficients $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ were all set to zero except at resonance conditions implied in Eqs.
+(4).
 Coupling coefficients were also supposed to be proportional to the square root of the masses and to the frequencies.
-For instance, in the case of $\phi$ coupling, the following coefficients were used:
-$\phi_{ik}=\left\{\begin{array}[]{l}\phi\,\sqrt{m_{i}m_{k}^{\scriptscriptstyle{(B)}}}\omega_{i}\omega_{k}^{\scriptscriptstyle(B)}/N_{\phi},\>\text{when}\ \omega_{i}=\omega_{k}^{\scriptscriptstyle(B)}\vspace{0.4cm}\\ 0\>\ \text{otherwise,}\end{array}\right.$ (5)
-where $\phi$ is a unitless parameter and $N_{\phi}$ is the number of resonances related to $\phi$ interactions.
-A similar expression was used for the $\xi_{il}$ coefficients by introducing the parameter $\xi$ and the number $N_{\xi}$ of $\xi$-resonances (not shown).
-Similarly, coefficients related to $\lambda$-coupling were set as
-$\lambda_{ijk}=\left\{\begin{array}[]{l}\lambda\,\sqrt{m_{i}m_{j}m_{k}^{\scriptscriptstyle{(B)}}}\omega_{i}\omega_{j}\omega_{k}^{\scriptscriptstyle(B)}/N_{\lambda},\vspace{0.1cm}\\ \qquad\qquad\qquad\text{when}\ \omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0\vspace{0.3cm}\\ 0\>\ \text{otherwise}\end{array}\right.$ (6)
-where $\lambda$ is given in $\text{K}^{-1/2}$ units.
-Assuming the bath and the source are much bigger entities than the protein, their frequencies spectra were always set to generate the maximum number of resonances possible, i.e., $N_{\phi}=N$, $N_{\xi}=N$ and $N_{\lambda}=N(N-1)$ where $N$ is the number of protein modes.
+For instance, in the case of $\phi$ coupling, the following coefficients were used when $\omega_{i}=\omega_{k}^{\scriptscriptstyle(B)}$:
+$\phi_{ik}=\phi\sqrt{m_{i}m_{k}^{\scriptscriptstyle{(B)}}}\omega_{i}\omega_{k}^{\scriptscriptstyle(B)}/N_{\phi},$ (5a)
+with $\phi_{ik}=0$ otherwise. Here $\phi$ is a unitless parameter and $N_{\phi}$ is the number of resonances related to $\phi$ interactions. A similar expression was used for $\xi_{il}$:
+$\indent\xi_{il}=\sqrt{m_{i}m_{l}^{\scriptscriptstyle{(S)}}}\omega_{i}\omega_{l}^{\scriptscriptstyle(S)}/N_{\xi}$ (5b)
+when $\omega_{i}=\omega_{l}^{\scriptscriptstyle(S)}$ ($\xi_{il}=0$ otherwise). Finally, coefficients related to $\lambda$-coupling were set as
+$\lambda_{ijk}=\lambda\sqrt{m_{i}m_{j}m_{k}^{\scriptscriptstyle{(B)}}}\omega_{i}\omega_{j}\omega_{k}^{\scriptscriptstyle(B)}/N_{\lambda}$ (5c)
+when $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ ($\lambda_{ijk}=0$ otherwise) with $\lambda$ given in $\text{K}^{-1/2}$ units.
+In all our simulations, the bath was composed of $500$ oscillators with frequencies uniformly distributed from $0.002$ to $1.0$ THz, with increments of $0.002$ THz, while the source consisted of $100$ oscillators uniformly distributed from $0.01$ to $1.0$ THz, with increments of $0.01$ THz.
 Physical intuition for setting $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ from Eqs.
-(5) and (6) is given in the next sections.
+(5) is given in the next sections.
 Results of our simulations are displayed in Fig.
-3, where we observed that the two main features of Fröhlich systems could be reproduced, that is, energy equipartition when nonlinear interactions are turned off ($\lambda=0$), and condensation in mode 1 at high $\lambda$ value.
-Although the FRE show reasonable agreement with real dynamics, they also tend to miscalculate the energy available in condensates, either by overestimating (Fig.
+3, where we reproduced the two main features of Fröhlich systems, that is, energy equipartition when nonlinear interactions are turned off ($\lambda=0$), and condensation in the lowest frequency mode at high $\lambda$ value.
+Although the FREs show reasonable agreement with real dynamics, they also tend to miscalculate the energy available in condensates, either by overestimating (Fig.
 3 bottom) or underestimating it (Fig.
-S3), showing the limits of Eqs.
+S3 in [15]), showing the limits of Eqs.
 (1) combined with Eq.
 (4).
-Dynamics also revealed the formation of strong condensates in specific regions of the parameter space characterized by reasonable source temperature and energies of a dozen of $kT$ (Fig.
-S3), consistent with typical energy values required to induce natural conformational changes in proteins.
-This observation goes against a previous numerical study by Reimers et al.
-affirming that strong condensates are unlikely to happen in a biological environment due to the high-energy values they require [11].
-Figure 4: Condensation index $\eta$ of a Fröhlich system as deduced from Hamiltonian dynamics for different $\lambda$ values.
-$\lambda_{ijk}$ coefficients were calculated from Eqs.
-(6) under different types of resonance: $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ (Fröhlich), $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$ (Lifshits) and $\omega_{i}\pm\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ (combination of both).
+Dynamics also revealed the formation of even more sizable condensates in other regions of the parameter space characterized by reasonable source and condensation energies (Fig.
+S3 in [15]).
+Energies in the range of $10^{3}-10^{4}$ K equivalent temperature actually correspond to about 2-20 kcal/mol, which is the typical order of magnitude of the energy barrier associated with functional conformational rearrangements in proteins.
+Interestingly, in light-activated proteins energy conversion occurring within the protein upon light illumination leads to local and global structural rearrangements [21, 22], which is compatible with the availability of a quite spatially delocalized and coherent form of energy.
+Phytochrome B photoreceptor exploits temperature to safely reverse from active to inactive conformational state [23], which proves how proteins can use a disordered and incoherent form of energy to accomplish a well-defined collective rearrangement.
+We note here that using the index and threshold value introduced in [8], the observed condensates would be classified as ”weak”.
+Based on the energetic argument above, the threshold given in [8] to discriminate biologically relevant/not relevant energies can be relaxed.
+Figure 4: Condensation index $\rho$ of a Fröhlich system as deduced from Hamiltonian dynamics for different $\lambda$ values.
+$\lambda_{ijk}$ coefficients were calculated from Eq.
+(5c) under different types of resonance: $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ (Fröhlich), $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$ (Lifshits) and $\omega_{i}\pm\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ (combination of both).
 Other parameters remain the same as Fig.
 3.
-Lifshits terms—We mentioned above that the FRE implicitly assume that nonlinear resonances of the type $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ are predominant over other types of resonance like $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$.
+Other coupling schemes— Our choice to apply Eq.
+(5c) only when $\omega_{i}-\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$ (Fröhlich resonances) was motivated by the fact that the FREs implicitly assume that such nonlinear resonances are predominant over other types of resonance like $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$.
 Lifshits [24] identified the latter type as a possible hindrance to condensation.
-To investigate this, another round of simulations was run by computing the $\lambda_{ijk}$ coefficients from the same expression given in Eq.
-(6) but setting them to be nonzero only when $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$.
+To investigate this, another round of simulations was run by computing the $\lambda_{ijk}$ coefficients from (5c) only when $\omega_{i}+\omega_{j}-\omega_{k}^{\scriptscriptstyle(B)}=0$ ($\lambda_{ijk}=0$ otherwise).
 Extra simulations were performed in the case $\omega_{i}\pm\omega_{j}\pm\omega_{k}^{\scriptscriptstyle(B)}=0$, i.e., when both Fröhlich and Lifshits resonances were involved.
 Results are depicted in Fig.
-4 highlighting the condensation index $\eta$ as a function of the $\lambda$ parameter.
-Here $\eta$ was defined as the energy of mode 1 in the stationary state divided by the average energy over the secondary modes, i.e., $\eta=E_{1}/\langle E_{i}\rangle_{i=2..N}$.
-Thus, $\eta=1$ indicates energy equipartition whereas $\eta>1$ implies condensation.
+4 highlighting the condensation ratio $\rho$ as a function of the $\lambda$ parameter.
+Here $\rho$ was defined as the energy of mode 1 in the stationary state divided by the average energy over the secondary modes, i.e., $\rho=E_{1}/\langle E_{i}\rangle_{i=2..N}$.
+Thus, $\rho=1$ indicates energy equipartition whereas $\rho>1$ implies condensation.
 From Fig.
@@
 Although it was suggested that the magnitude of Lifshits resonances is likely to be negligible in practice [18], investigation of Fröhlich condensates in real structures would require careful evaluation of these contributions as deduced from all-atom force fields.
+Moving further away from the ideal Fröhlich resonance conditions, we also investigated the regime in which the resonance constraints on both linear and nonlinear coupling terms are fully relaxed.
+In this scenario, $\phi_{ik}$ , $\xi_{il}$ and $\lambda_{ijk}$ were set from Eqs.
+(5a), (5b) and (5c), respectively, for all $i,k$ and $l$.
+In this case, no condensate forms in the dynamics, possibly owing to Lifshits resonances that could interfere with the condensation mechanism.
+Since Lifshits resonances typically involve higher bath frequencies than Fröhlich ones, nonlinear interactions with a low-pass bath were tested.
+Specifically, the $\lambda_{ijk}$’s were computed from (5c) only when $\omega_{k}^{\scriptscriptstyle(B)}\leq\omega_{0}^{\scriptscriptstyle(B)}$ ($\lambda_{ijk}=0$ otherwise).
+Here $\omega_{0}^{\scriptscriptstyle(B)}$ refers to a typical bath frequency beyond which Lifshits resonances become dominant over Fröhlich ones (Fig.
+S4 in [15]).
+This type of nonlinear coupling was able to restore energy condensation (Fig.
+S5 in [15]), suggesting that low bath frequencies are crucial for redistributing energy into the lowest-frequency protein mode (see also Table 1 for a summary of conditions leading to condensation).
+Table 1: Resonance conditions leading to condensation; $\lambda_{ijk}$ is computed from Eq.
+(5c) when the condition is satisfied and set to 0 otherwise.
+In the first 3 cases, linear coefficients $\phi_{ik}$ and $\xi_{il}$ are computed from Eqs.
+(5a) and (5b) only when $\omega_{i}=\omega_{k}^{\scriptscriptstyle(B)}$ and $\omega_{i}=\omega_{l}^{\scriptscriptstyle(S)}$ and set to 0 otherwise.
+For the last 2 cases, linear coefficients are computed from the same equations for all $i$, $k$, and $l$.
+Case Condition on $\lambda_{ijk}$ Condensation
+Fröhlich $\omega_{i}-\omega_{j}\pm\omega_{k}^{(B)}=0$ $\checkmark$
+Lifshits $\omega_{i}+\omega_{j}-\omega_{k}^{(B)}=0$ $\times$
+All resonances $\omega_{i}\pm\omega_{j}\pm\omega_{k}^{(B)}=0$ $\times$
+Low-pass bath $\omega_{k}^{\scriptscriptstyle(B)}\leq\omega_{0}^{\scriptscriptstyle(B)}$ $\checkmark$
+General no restrictions $\times$
 Protein-bath vs protein-protein coupling—In addition to nonlinear resonances, we explored the importance of the heat bath in nonlinear coupling.
@@
 While we cannot rule out the existence of Fröhlich condensates in unexplored regions of the parameter space, our results strongly indicate that bath-mediated coupling is essential for inducing condensation.
-Constant coupling—In their study, Reimers et al.
-[11] reported that they had investigated Fröhlich condensation over a wide region of the parameter space of the quantum Wu-Austin Hamiltonian, concluding with the non-existence of strong and coherent condensates in living matter.
-However, only constant coupling parameters were considered in their Hamiltonian, drastically limiting the exploration of the condensates.
-As mentioned above, our simulations were run by setting $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ proportional to the square root of the masses and to the frequencies.
+Constant coupling —Our simulations were run by setting $\phi_{ik}$, $\xi_{il}$ and $\lambda_{ijk}$ proportional to the square root of the masses and to the frequencies.
 Such a choice was motivated by the fact that, at thermal equilibrium, averaged potential energies $\frac{1}{2}m_{i}\omega_{i}^{2}\langle q_{i}^{2}\rangle$ should equal $kT/2$ or $kT_{S}/2$ for the bath or the energy source, respectively.
-Thus, 2 modes with the same mass but different frequencies, say $\omega_{1}<\omega_{2}$, should satisfy $\langle q_{1}^{2}\rangle>\langle q_{2}^{2}\rangle$ in order to preserve equilibrium condition, meaning that low-frequency modes will have naturally higher amplitude than high-frequency ones.
+Thus, 2 modes with the same mass but different frequencies, say $\omega_{1}<\omega_{2}$, should satisfy $\langle q_{1}^{2}\rangle>\langle q_{2}^{2}\rangle$ in order to preserve equilibrium condition, meaning that low-frequency modes will naturally have higher amplitude than high-frequency ones.
 In that case, an interaction potential with constant coefficients (e.g., $\phi_{ik}q_{i}q_{k}^{{\scriptscriptstyle(B)}}$ with $\phi_{ik}=\phi$) will generate stronger interactions between low-frequency modes.
 This observation can also be made for masses, i.e., modes with a large mass will have a stronger impact on the interaction energy.
-Setting coefficients as in Eqs (5) and (6) enables to circumvent this issue, ensuring that all the modes contribute equally to the potential.
+Setting coefficients as in Eqs (5) enables to circumvent this issue, ensuring that all the modes contribute equally to the potential.
 Despite the above, the impact of a constant coupling on a Fröhlich system was still investigated.
@@
 This result is clearly not compatible with our Hamiltonian model whereby $H_{int}$ is always considered a small perturbation as compared to $H_{0}$.
-Again, using Eqs (5) and (6) helped fix this issue by keeping the interaction energy reasonably low even when strong condensation occurs (Fig.
-S4, see also note [20]).
+Again, using Eqs (5) helped fix this issue by keeping the interaction energy reasonably low even when sizable condensation occurs (Fig.
+S6 in [15]).
+Thermostats—Aside from nonlinear terms and the choice of coupling coefficients, thermostats used to maintain the bath and the energy source at the target temperature may also have a non-negligible impact on the emergence of condensation.
+The influence of the Langevin friction coefficient of the bath ($\gamma$) and the source ($\gamma_{S}$) was investigated by varying each individually (Table S1) or both simultaneously (Table S2) from 0.03 to 1 ps-1, while keeping other parameters fixed.
+Results show that condensation is overall robust and its extent can be modulated by varying $\gamma$ and/or $\gamma_{S}$, highlighting the importance of dissipation and injection rates on Fröhlich effect.
+Rescaling study—To further assess the robustness of the condensation mechanism and the impact of the parameters on the timescale of the condensation, coupling coefficients were all rescaled by the same multiplicative factor $\kappa$ over two orders of magnitude.
+Parameters used to produce Fig.
+3 were taken as a reference ($\kappa=1.0$).
+The $\kappa$ factor was then varied from $0.01$ to $1$ and applied simultaneously to $\phi$, $\xi$, and $\lambda$, while the friction coefficients $\gamma$ and $\gamma_{S}$ are left unchanged.
+Fig.
+S7 in [15] shows the condensation ratio $\rho$ as a function of $\kappa$, along with the corresponding condensation time.
+The latter was obtained by fitting the average energy of the lowest-frequency mode as a function of time.
+Fitting was performed using the function $A-B\exp{(-t/\tau)}$ from the moment the source is activated ($\xi\neq 0$) when the energy is at the equipartition value of 300 K.
+The condensation time was then defined as the fitted coefficient $\tau$ multiplied by $\ln(2)$ corresponding to the time to reach half of the plateau value.
+From our analysis, we see that $\rho$ strongly increases as $\kappa$ decreases, while the condensation time follows a similar trend but rises much more sharply at small $\kappa$ values.
+Figure 5: Blue: condensation ratio $\rho$ as a function of the scaling factor $\kappa$ multiplying the coupling coefficients $\phi$, $\xi$, and $\lambda$ and the friction coefficients $\gamma$ and $\gamma_{S}$ simultaneously.
+$\kappa=1.0$ corresponds to the parameters used in Fig.
+3.
+Except for $\phi$, $\xi$, $\lambda$, $\gamma$, and $\gamma_{S}$, all other parameters are the same as in Fig.
+3 for all $\kappa$ values.
+Green: condensation time (see main text for details).
+The same rescaling study was repeated by applying the $\kappa$ factor to both the friction coefficients $\gamma$ and $\gamma_{S}$, and to the coupling constants.
+Results are displayed in Fig.
+5.
+As in the previous case, a steep increase of the condensation time is observed as $\kappa$ becomes smaller.
+Although $\rho$ also increases, a clear concave dependence is found in this case, with $\rho$ stabilizing at very small $\kappa$.
+This highlights conditions in which the condensation time can grow to very large values while keeping the condensed energy relatively stable.
 Discussion and perspectives—The present study deals with the foundations of the condensation phenomenon proposed by Fröhlich in the 1960s aimed at explaining coherent excitations in biomolecular systems [5].
-Motivated by recent experimental findings [2, 3], we simulated the dynamics of a general classical Hamiltonian that incorporates all the key elements of Fröhlich’s models.
-Our results show that condensation, including strong condensates with reasonable energy values, can emerge under specific types of nonlinear coupling, supporting the physical plausibility of Fröhlich’s effect.
-However, small deviations from the original model, for example, via the addition of resonances involving high-frequency modes of the bath or the use of a nonlinear potential without bath mediation, result in suppressing the phenomenon.
+Motivated by recent experimental findings [2, 3, 4], we simulated the dynamics of a general classical Hamiltonian that incorporates all the key elements of Fröhlich’s models.
+Our results show that condensation, including strong condensates with reasonable energy values, can emerge under specific types of nonlinear coupling (Fröhlich resonances), supporting the physical plausibility of Fröhlich’s effect.
+Conversely, the addition of resonances involving high-frequency modes of the bath (Lifshits resonances) or the use of a nonlinear potential without bath mediation, result in suppressing the phenomenon.
+Interestingly, condensation still emerges without specific resonance conditions provided that nonlinear interactions are driven by low-frequency bath modes.
+This was implemented by introducing a low-pass filter on the bath frequencies, where the threshold is the frequency at which the number of verified Lifshits resonances exceeds that of Fröhlich’s.
 These factors should be carefully examined when investigating the condensation effect in real biomolecular structures.
 Given the large number of free parameters involved in the dynamics of Fröhlich’s systems, a full exploration of the parameter space is beyond the scope of this letter.
-Although off-resonance coupling [21] or higher-order coupling mechanisms could bring more realistic characteristics to the model [22], a more practical approach would be to tune coupling parameters directly from standard MD force fields.
+Although higher-order coupling mechanisms could bring more realistic characteristics to the model [25], a more practical approach would be to tune the order-three coupling parameters so far considered directly from standard MD force fields targeting a specific model protein.
 Possible strategies include normal mode analysis (NMA) to generate the set of eigenvectors needed to transition from real space to the space of protein normal modes [26, 27].
@@
 (3) to be deduced and used in our model to verify whether condensation can indeed arise.
-Finally, such an approach should give further insights into the reasons why previous MD studies have been unsuccessful in detecting Fröhlich’s effect and if missing ingredients are required.
+Furthermore, we note that energy condensation remains sizable for a quite large interval of source and bath friction coefficients, including typical values used in MD simulations.
+Finally, while our frequencies are representative of low-frequency modes in proteins, coupling constants values were essentially chosen to observe a well-formed condensate within a reasonable timescale, typically on the order of several hundreds of ns.
+Reducing the magnitude of the coupling constants in our Hamiltonian was found to delay the formation of condensates.
+Although at this stage one cannot draw a direct link between the timescales explored by a typical MD force field and those explored in our model, our findings suggest that the length of force field-based simulations could be a possible limiting factor in observing a clear signature of well-formed Fröhlich’s condensates.
 Acknowledgments—This work and J.